# Relatives of the Hermitian curve

### v2

@@ 3 Example I - diagonal case
$\begin{array}[]{rcccccccl}0\to&\mathbb{F}_{\sqrt{q}}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}&\to&\mathbb{F}_{q}&\xrightarrow{\operatorname{Tr}}&\mathbb{F}_{\sqrt{q}}&\to 0\\ &&&\alpha&\mapsto&\alpha^{\sqrt{q}}-\alpha&&&\end{array};$ (5)
$\begin{array}[]{rcccccccl}1\to&\mathbb{F}_{\sqrt{q}}^{\ast}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}^{\ast}&\to&\mathbb{F}_{q}^{\ast}&\xrightarrow{\operatorname{Nm}}&\mathbb{F}_{\sqrt{q}}^{\ast}&\to 1\\ &&&\alpha&\mapsto&\alpha^{\sqrt{q}-1}&&&\end{array},$ (6)
@@
Proof.
This lemma is a special case of Hilbert’s Theorem 90, and the proof is easy. ∎
Corollary 3.2
@@
(i) Let $\alpha\in\mathbb{F}_{q}$ be a root of (7). Then $\beta=\alpha^{\sqrt{q}}-\alpha$. Hence $\operatorname{Tr}\beta=0$ by (5). Conversely, if $\operatorname{Tr}\beta=0$, there is an element $\alpha\in\mathbb{F}_{q}$ such that $\alpha^{\sqrt{q}}-\alpha=\beta$ by (5) again. When $\alpha$ is a root of (7), the set of all roots is $\{\alpha+\gamma|\gamma\in\mathbb{F}_{\sqrt{q}}\}$.
@@
In this section, we handle only diagonal matrices; $\operatorname{diag}[a,b,c]$ denotes the diagonal matrix whose diagonal entries are $a,b,c$ from the upper left to the lower right.
Let $A\in PGL(3,\mathbb{F}_{q})$, which is diagonal. Then $A$ is equivalent to one of the following three types.
Example 3.3
(Type 0) $A\sim\operatorname{diag}[1,1,1]$. So the corresponding curve $C_{A}$ is projectively equivalent to the Hermitian curve $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0$, and $N_{q}(C_{A})=(\sqrt{q})^{3}+1$. Hence $m_{C_{A}}=q$.
Example 3.4
(Type 1) $A\sim\operatorname{diag}[1,1,\omega]$ for some $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. Hence $C_{A}$ is projectively equivalent to the curve $D_{\omega}$ defined by
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\omega z^{\sqrt{q}+1}=0.$ (9)
@@
because $x_{0}^{\sqrt{q}+1},y_{0}^{\sqrt{q}+1},z_{0}^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}$ and $\{1,\omega\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$. Therefore
$D_{\omega}(\mathbb{F}_{q})=\{(\zeta,1,0)|\operatorname{Nm}\zeta=-1\}.$
Hence $N_{q}(C_{A})=\sqrt{q}+1$, and $m_{C_{A}}=1$.
Example 3.5
(Type 2) $A\sim\operatorname{diag}[1,\omega,\omega^{\prime}]$ for some $\omega,\omega^{\prime}\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$ with $\omega(\omega^{\prime})^{-1}\not\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Hence $C_{A}$ is projectively equivalent the curve $D_{\omega,\omega^{\prime}}$ defined by
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+\omega^{\prime}z^{\sqrt{q}+1}=0.$ (10)
Since any two elements of $\{1,\omega,\omega^{\prime}\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, there is a unique relation up to $\mathbb{F}_{\sqrt{q}}^{\ast}$-multiple among those three elements, say
$\alpha+\beta\omega+\gamma\omega^{\prime}=0.$
Since none of three coefficients $\alpha,\beta,\gamma$ are 00, we may assume that $\alpha=1$, and the first coordinate of any solution of (10) is also $1$. Then $D_{\omega,\omega^{\prime}}(\mathbb{F}_{q})=\{(1,\zeta,\eta)|\operatorname{Nm}\zeta=\beta,\operatorname{Nm}\eta=\gamma\}$. Hence $N_{q}(C_{A})=(\sqrt{q}+1)^{2}$, and hence $m_{C}=\sqrt{q}+2$.
4 Example II - non-diagonal case
The following examples, including the diagonal case explained above, are inspired by Hirschfeld’s classification of cubic plane curves over any finite fields [2, Section 11]. Actually, the family of supersingular plane cubic curves over $\mathbb{F}_{4}$ coincides with the family of relatives of the Hermitian curve over $\mathbb{F}_{4}$, because any curves in each family is characterized by disappearance of the term $xyz$ in its equation (see, [1, IV, Prop. 4.21]).
Example 4.1
Let $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$, and
$A=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}.$
Then $C_{A}$ is defined by
$x^{\sqrt{q}}y+xy^{\sqrt{q}}+\omega z^{\sqrt{q}+1}=0.$
The curve $C_{A}$ is projectively equivalent over $\mathbb{F}_{q}$ to a diagonal curve of type 1. In fact, choose $u\in\mathbb{F}_{q}$ so that $\operatorname{Tr}u=1$ and $\alpha\in\mathbb{F}_{q}^{\ast}$ so that $\operatorname{Nm}\alpha=-1$. Let $T_{0}=\begin{pmatrix}u&\alpha u^{\sqrt{q}}&0\\ 1&-\alpha&0\\ 0&0&1\end{pmatrix}.$ Then it is easy to see that
${}^{t}\!T_{0}^{(\sqrt{q})}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}T_{0}=\operatorname{diag}[1,1,\omega].$
Example 4.2
Let us consider the matrix $A=\begin{pmatrix}1&0&0\\ 0&\omega&1\\ 0&1&0\end{pmatrix},$ where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. So the curve $C_{A}$ is defined by
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$ (11)
Proposition 4.3
For the curve (11), $C_{A}(\mathbb{F}_{q})=\{(0,0,1)\}$, and hence $m_{C_{A}}=0$.

### v3

@@ 3.1 Some equations
$\begin{array}[]{rcccccccl}0\to&\mathbb{F}_{\sqrt{q}}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}&\to&\mathbb{F}_{q}&\xrightarrow{\operatorname{Tr}}&\mathbb{F}_{\sqrt{q}}&\to 0\\ &&&\alpha&\mapsto&\alpha-\alpha^{\sqrt{q}}&&&\end{array}$ (7)
$\begin{array}[]{rcccccccl}1\to&\mathbb{F}_{\sqrt{q}}^{\ast}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}^{\ast}&\to&\mathbb{F}_{q}^{\ast}&\xrightarrow{\operatorname{Nm}}&\mathbb{F}_{\sqrt{q}}^{\ast}&\to 1\\ &&&\alpha&\mapsto&\alpha^{1-\sqrt{q}}&&&\end{array},$ (8)
@@
Corollary 3.3.
@@
(i) Let $\alpha\in\mathbb{F}_{q}$ be a root of (9). Then $\beta=\alpha^{\sqrt{q}}-\alpha$. Hence $\operatorname{Tr}\beta=0$. Conversely, if $\operatorname{Tr}\beta=0$, there is an element $\alpha\in\mathbb{F}_{q}$ such that $\alpha^{\sqrt{q}}-\alpha=\beta$ by (7). When $\alpha$ is a root of (9), the set of all roots is $\{\alpha+\gamma|\gamma\in\mathbb{F}_{\sqrt{q}}\}$.
@@
The above corollary can be generalized slightly as follows:
Proposition 3.4.
Let $\alpha\in\mathbb{F}_{q}^{\ast}$ and $\beta\in\mathbb{F}_{q}$, and consider the equation in $X$:
$X^{\sqrt{q}}+\alpha X+\beta=0.$ (11)
  1. (i)
[…]
  2. (ii)
[…]
Proof.
For a fixed $\alpha\in\mathbb{F}_{q}^{\ast}$, consider the $\mathbb{F}_{\sqrt{q}}$-linear map $\varphi_{\alpha}:\mathbb{F}_{q}\to\mathbb{F}_{q}$ given by $\varphi_{\alpha}(z)=z^{\sqrt{q}}+\alpha z$. Since $\operatorname{Nm}(-\alpha)=\operatorname{Nm}\alpha$, the kernel of $\varphi_{\alpha}$ is nontrivial if and only if $\operatorname{Nm}\alpha=1$ by Corollary 3.3 (ii).
Therefore if $\operatorname{Nm}\alpha\neq 1$, then $\varphi_{\alpha}$ is an isomorphism. Hence there is a unique element of $\mathbb{F}_{q}$, say $z_{0}$, so that $\varphi_{\alpha}(z_{0})=-\beta$. So the proof of (ii) is done.
Next suppose that $\operatorname{Nm}\alpha=1$. In this case, $\dim\operatorname{Ker}\varphi_{\alpha}>0$ from the first paragraph of this proof, in particular there is an element $z_{0}\in\mathbb{F}_{q}^{\ast}$ such that $z_{0}^{\sqrt{q}-1}+\alpha=0$. Hence
$\operatorname{Ker}\varphi_{\alpha}=\{\zeta z_{0}\mid\zeta\in\mathbb{F}_{\sqrt{q}}^{\ast}\}\cup\{0\}.$
If (11) has a solution $z_{1}$, the set of all solutions of (11) is
$\{z_{1}+\zeta z_{0}\mid\zeta\in\mathbb{F}_{\sqrt{q}}^{\ast}\}\cup\{z_{1}\},$
which is a subset of $\mathbb{F}_{q}$. So the additional statement of (i) is done.
Now we prove the first part of (i). To prove the ‘only if’ part, we may suppose that $\beta\neq 0$. Let $z_{1}$ be an $\mathbb{F}_{q}$-solution of (11). Then $\alpha=-\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}}$. Since $\operatorname{Nm}\alpha=1$ and $(z_{1}^{\sqrt{q}})^{\sqrt{q}}=z_{1}$, we have
11 11 $=$ $\operatorname{Nm}\alpha=-(\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}})^{\sqrt{q}}(-\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}})$
$=$ $\frac{z_{1}^{\sqrt{q}+1}+\beta^{\sqrt{q}}z_{1}^{\sqrt{q}}+\beta z_{1}+\beta^{\sqrt{q}+1}}{z_{1}^{\sqrt{q}+1}}.$
Hence
$z_{1}^{\sqrt{q}}+\beta^{1-\sqrt{q}}z_{1}+\beta=0.$
Since $z_{1}$ satisfies (11), we get $\alpha=\beta^{1-\sqrt{q}}$. Conversely, if $\beta=0$, obviously (11) has an $\mathbb{F}_{q}$-solution. Hence we may consider the case $\alpha=\beta^{1-\sqrt{q}}$ with $\beta\neq 0$. In this case, (11) is
$X^{\sqrt{q}}+\beta^{1-\sqrt{q}}X+\beta=0$ (12)
which is equivalent to
$(\beta X)^{\sqrt{q}}+(\beta X)+\beta^{\sqrt{q}+1}=0.$ (13)
Since
$(-\beta^{\sqrt{q}+1})^{\sqrt{q}-1}=\beta^{q-1}=1$
because $\beta\in\mathbb{F}_{q}^{\ast}$, we know $-\beta^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Since $\operatorname{Tr}:\mathbb{F}_{q}\to\mathbb{F}_{\sqrt{q}}$ is surjective, the equation (13) in $(\beta X)$ has a solution in $\mathbb{F}_{q}$, and also (12) has. ∎
3.2 Proof of Theorem 3.1
Lemma 3.5.
Let $C$ be a relative of the Hermitian curve with at least two rational inflexions. Then we can choose a system of coordinates $x,y,z$ of $\mathbb{P}^{2}$ over $\mathbb{F}_{q}$ such that
  • $P=(1,0,0)$ and $Q=(0,1,0)$ are inflexions of $C$, and
  • $T_{P}(C)=\{y=0\}$ and $T_{Q}(C)=\{x=0\}$.
Proof.
Since $C.T_{P}(C)=(\sqrt{q}+1)P$ and $C.T_{Q}(C)=(\sqrt{q}+1)Q$, three $\mathbb{F}_{q}$-lines $T_{P}(C)$, $T_{Q}(C)$ and $PQ$ are not concurrent, where $PQ$ denotes the line through $P$ and $Q$. Therefore, we may choose coordinates as $T_{P}(C)=\{y=0\}$, $T_{Q}(C)=\{x=0\}$ and $PQ=\{z=0\}$. ∎
Lemma 3.6.
Under the situation described in Lemma 3.5, the curve is defined by $(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\,{}^{t}\!(x,y,z)=0,$ where
$A=\begin{pmatrix}0&a_{12}&0\\ a_{21}&0&0\\ 0&0&a_{33}\end{pmatrix}\in PGL(3,\mathbb{F}_{q})\ \text{with}\ a_{12}a_{21}a_{33}\neq 0.$
In other words, $C$ is defined by
$a_{12}x^{\sqrt{q}}y+a_{21}xy^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0.$ (14)
Proof.
Let $C=C_{A}$, where $A=(a_{ij})\in PGL(3,\mathbb{F}_{q})$. Since $(1,0,0)A{}^{t}\!(1,0,0)=0$ and $(0,1,0)A{}^{t}\!(0,1,0)=0$, $a_{11}=a_{22}=0$.
Note that the tangent line at $(b_{1},b_{2},b_{3})\in C$ is given by
$(b_{1}^{\sqrt{q}},b_{2}^{\sqrt{q}},b_{3}^{\sqrt{q}})A{}^{t}\!(x,y,z)=0.$
Hence $a_{13}=0$ because $T_{P}(C)=\{y=0\}$, and $a_{23}=0$ because $T_{Q}(C)=\{x=0\}$. Therefore an equation of $C$ is of the form
$a_{12}x^{\sqrt{q}}y+a_{21}xy^{\sqrt{q}}+a_{31}xz^{\sqrt{q}}+a_{32}yz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0.$
Since $P$ is an inflexion of $C$ and $T_{P}(C)=\{y=0\}$, we have $C.\{y=0\}=(\sqrt{q}+1)P$. Hence the equation
$a_{31}xz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0$
has the root at $z=0$ of multiplicity $(\sqrt{q}+1)$. So $a_{31}=0$. Similarly, since $C.\{x=0\}=(\sqrt{q}+1)Q$, from the equation $a_{32}yz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0,$ we know $a_{32}=0$. Finally $\det A\neq 0$ implies $a_{12}a_{21}a_{33}\neq 0$. ∎
Corollary 3.7.
After a suitable projective transformation over $\mathbb{F}_{q}$, $C$ is defined by
$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0$
with $\omega\in\mathbb{F}_{q}^{\ast}$. Conversely, a curve defined by this type of equation is Hermitian-relative with two or more inflexions.
Proof.
Since $a_{33}\neq 0$, we may suppose that $a_{33}=1$ in (14). If one consider $a_{12}y$ as new $y$, then the equation becomes
$x^{\sqrt{q}}y+a_{21}a_{12}^{-{\sqrt{q}}}xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0.$
Put $a_{21}a_{12}^{-{\sqrt{q}}}=\omega.$
To confirm the last assertion is not difficult, actually $(1,0,0)$ and $(0,1,0)$ are inflexions. ∎
By Corollary 3.7, we have done the proof of the first paragraph of Theorem 3.1.
Now we investigate the curve (5), hereafter, this curve will be denoted by $C_{\omega}$.
Lemma 3.8.
The number of $\mathbb{F}_{q}$-points on $C_{\omega}\cap\{z=0\}$ is $\sqrt{q}+1$ if $\operatorname{Nm}\omega=1$, and $2$ if $\operatorname{Nm}\omega\neq 1$.
Proof.
For any $\omega$, $P=(1,0,0)$ and $Q=(0,1,0)$ are rational inflexions on $C_{\omega}$. Since $\{z=0\}\cap\{y=0\}=\{P\}$, it is enough to count the number of $\mathbb{F}_{q}$-solutions of (5) when $y=1$ and $z=0$. The equation in question is $x^{\sqrt{q}}+\omega x=0.$ Except for $Q$, the other points of $C_{\omega}\cap\{z=0\}\cap\{y=1\}$ correspond to the roots of $x^{\sqrt{q}-1}+\omega=0.$ By Corollary 3.3, the number of $\mathbb{F}_{q}$-solutions of this equation is
$\left\{\begin{array}[]{cl}\sqrt{q}-1&\text{if $\operatorname{Nm}\omega\,(=\operatorname{Nm}(-\omega))=1$}\\ 0&\text{ else.}\end{array}\right.$
Adding $2$ coming from $\{P,Q\}$ in each cases, we have the desired numbers. ∎
Next we compute the number of rational points of $C_{\omega}$ outside of $\{z=0\}$.
Lemma 3.9.
The number of $\mathbb{F}_{q}$-solutions of affine equation
$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+1=0$ (15)
is
$\left\{\begin{array}[]{cl}\sqrt{q}(q-1)&\text{ if $\omega=1$}\\ 0&\text{ if $\operatorname{Nm}\omega=1$ and $\omega\neq 1$}\\ q-1&\text{ if $\operatorname{Nm}\omega\neq 1$}.\end{array}\right.$
Proof.
If the affine equation (15) has an $\mathbb{F}_{q}$-solution $(x_{0},y_{0})$, then $x_{0}y_{0}^{\sqrt{q}}$ is a root of the equation in $T$
$T^{\sqrt{q}}+\omega T+1=0.$ (16)
Conversely if $t_{0}$ is a root of (16), then $(\frac{t_{0}}{y_{0}^{\sqrt{q}}},y_{0})$ is a solution of (15) for any $y_{0}\in\mathbb{F}_{q}^{\ast}$. Therefore the number of solutions of (15) is
[…]
Suppose that $\operatorname{Nm}\omega=1$. Then, by Proposition 3.4 (i), (16) has an $\mathbb{F}_{q}$-solution if and only if $\omega=1$. In this case, it has exactly $\sqrt{q}$ rational solutions, and totally (15) has $\sqrt{q}(q-1)$ rational solutions.
Suppose that $\operatorname{Nm}\omega\neq 1$. Then, from the second part of the same proposition, (16) has a unique $\mathbb{F}_{q}$-solution, and then (15) has $(q-1)$ rational solutions. ∎
The following three lemmas will be used for showing the part (b) of Theorem 3.1.
@@
The element of $PGL(3,\mathbb{F}_{q})$ which comes from the diagonal matrix with entries $a,b,c$ from the upper left to the lower right will be denoted by $\operatorname{diag}[a,b.c]$.
Lemma 3.10.
For $\eta\in\mathbb{F}_{q}^{\ast}$, we have
$\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\eta\end{pmatrix}\sim\operatorname{diag}[1,1,\eta].$
Proof.
Choose $u\in\mathbb{F}_{q}$ so that $\operatorname{Tr}u=1$ and $\alpha\in\mathbb{F}_{q}^{\ast}$ so that $\operatorname{Nm}\alpha=-1$. Let $T_{0}=\begin{pmatrix}u&\alpha u^{\sqrt{q}}&0\\ 1&-\alpha&0\\ 0&0&1\end{pmatrix}.$ Then by straightforward computation, we know that $T_{0}^{\ast}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}T_{0}=\operatorname{diag}[1,1,\eta].$
Lemma 3.11.
Let $\eta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$. Then the number of rational points of $C_{\operatorname{diag}[1,1,\eta]}$ is $\sqrt{q}+1$, and these rational points are on the line $\{z=0\}$.
Proof.
The curve $C_{\operatorname{diag}[1,1,\eta]}$ is defined by
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0.$ (17)
@@
because $x_{0}^{\sqrt{q}+1},y_{0}^{\sqrt{q}+1},z_{0}^{\sqrt{q}+1}$ are elements of $\mathbb{F}_{\sqrt{q}}$, and $\{1,\eta\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$. Therefore
$C_{\operatorname{diag}[1,1,\eta]}(\mathbb{F}_{q})=\{(\zeta,1,0)|\operatorname{Nm}\zeta=-1\}\subset\{z=0\}.$
This completes the proof. ∎
Lemma 3.12.
Let $\eta,\eta^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$. Then there is an element $T\in PGL(3,\mathbb{F}_{q})$ such that
$T^{\ast}\operatorname{diag}[1,1,\eta]T=\operatorname{diag}[1,1,\eta^{\prime}]$
if and only if $\eta^{\prime}\eta^{-1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$.